OpenStudy (anonymous):

PLS HELP. WILL GIVE MEDALS> PLS ANSWer ALL solve each equation on interval {0,2pi} 1. 4sin^2x-3=0 a) rewrite so its expressed in terms of sin^2x=a. b)solve for sinx by taking the square root. include all possible values that satisfy the equation c) solve each equation separately to get the full solution set. choose only solutions in the interval. 2. cos(3x)=-1 a) write an equation that defines the value of 3x, inverse trig expression. b)solve to find all possible values of angle 3x c) use algreba to find values of x between 0 and 2pi to satisfy the equation.

OpenStudy (danjs):

what did you do so far

OpenStudy (anonymous):

nothing i dont understand it please go through all the steps and solve thanks

OpenStudy (danjs):

can you do the algebra to do a and b

OpenStudy (danjs):

Solve \[4*[\sin(x)]^2-3 = 0\] for sin(x)

jhonyy9 (jhonyy9):

so than derange you sin(x) than sign it sin(x)=a so than will get 4a^2 -3 = 0 and solve it for a 4a^2 = 3 a^2 = 3/4 a_1,2 = +/- sqrt(3/4) = +/- (sqrt3)/2 so go back and make it sin(x) = +/- (sqrt3)/2 can you ending it now ? - for this one little help - you need to know that sin60° = (sqrt3)/2 hope so much that these will help you